### v1

# COSINE-100U: Upgrading the COSINE-100 Experiment for Enhanced Sensitivity to Low-Mass Dark Matter Detection

Abstract
[…]
I Introduction
Numerous astronomical observations suggest that the majority of matter in the universe consists of invisible dark matter, though its nature and interactions remain elusive Clowe et al. (2006); Aghanim et al. (2020); Bertone and Hooper (2018). Despite extensive efforts to directly detect dark matter, no definitive signals have been observed Marrodán Undagoitia and Rauch (2016); Schumann (2019). The only exception is the DAMA/LIBRA experiment, which has reported annual modulation signals in a 250 kg array of NaI(Tl) detectors Bernabei et al. (1998, 2013, 2018, 2021), indicating potential dark matter-nuclei interactions Savage et al. (2009); Ko et al. (2019). However, these results remain controversial, as no other experiments have observed similar signals Schumann (2019); Workman et al. (2022), necessitating independent verification using the same NaI(Tl) crystal target materials.
[…]
The COSINE-100 experiment concluded in March 2023 after 6.4 years of stable operation. It was relocated to the newly constructed Yemilab Park et al. (2024); Kim and Lee (2024) in Korea for the next phase, COSINE-100U. During the transition, the crystal encapsulation was upgraded to improve light collection by directly attaching the crystals to photomultiplier tubes (PMTs) Choi et al. (2020). A similar technique, applied in the NEON experiment for measuring coherent neutrino-nucleus scattering at nuclear reactors Choi et al. (2023), achieved approximately a 50% improvement in light yield and demonstrated long-term stability over two years Choi et al. (2024a). These improvements in light yield are expected to lower the energy threshold, thereby improving the sensitivity of COSINE-100U to dark matter, particularly in the low-mass region. For the COSINE-100U experiment at Yemilab, all eight NaI(Tl) crystals from COSINE-100 were upgraded. In this article, we detail the detector upgrade and the resulting improved sensitivities for low-mass dark matter detection.
II NaI(Tl) Crystal Encapsulation
Commercial companies typically package NaI(Tl) crystals using aluminum or copper bodies with quartz windows. The crystal surfaces are often wrapped in reflective materials, such as Teflon sheets or aluminum oxide powder, to enhance light reflection.
All eight COSINE-100 crystals were grown by Alpha Spectra Inc. in collaboration with low-background research efforts of the KIMS, DM-Ice, and ANAIS collaborations Kim et al. (2015); Amaré et al. (2015); Adhikari et al. (2016). These cylindrical crystals were hermetically encased in oxygen-free copper (OFC) tubes (1.5 mm thick) with quartz windows at both ends, as exemplified by crystal-6 (C6) in Fig. 1. The lateral surfaces of each crystal were wrapped in Teflon sheets before being inserted into the OFC tubes. A 12 mm-thick quartz window was coupled to the crystal using a 1.5 mm-thick optical pad, and PMTs were attached to the quartz window using a small amount of high-viscosity optical grease.
Figure 1: The encapsulation design of the COSINE-100 C6 crystal. A 4.8-inch diameter crystal is encased in OFC tubes, with an optical pad and quartz windows. The 12 mm-thick quartz window guides light from the 4.8-inch diameter crystal to the 3-inch PMTs, aided by a 45 angle surface. There are three optical interface layers between the crystal and the PMTs: a 1.5 mm optical pad, a 12 mm quartz windows, and a small amount of optical grease. The PMTs are encapsulated in OFC cases to protect them from the surrounding liquid scintillator.
From the perspective of light collection efficiency, the original COSINE-100 encapsulation design had several drawbacks. The three-layer optical interface–comprising the optical pad, quartz window, and optical grease–added additional reflections, reducing light collection efficiency. Although the 12 mm-thick quartz window, with a 45 angle, guided photons from the 4.8-inch diameter crystal, it was insufficient for efficiently directing photons to the 3-inch PMTs. As shown in Fig. 1, uncovered areas resulted in photon reflection, further reducing light collection efficiency.
To address these issues, we developed a novel crystal encapsulation technique that involved directly attaching the PMT to the crystal using only a 2 mm-thick optical pad Choi et al. (2020). In this method, the crystal’s diameter was matched to the 3-inch diameter of the PMTs, eliminated photon absorption in the quartz and minimized photon loss due to reflections from multiple interfaces. This modification resulted in up to 22 photoelectrons (NPE)/keV, approximately 50% higher than the light yield with the COSINE-100 crystals Adhikari et al. (2018a).
In this design, the crystal and PMTs were assembled as a single unit, with the PMTs becoming integral components of the airtight crystal encapsulation. A 1 mm-thick PTFE o-ring was used to seal the neck of the PMT glass between the copper cylinder’s endcap and the back of the PMT glass, preventing any air leakage. This technique was initially applied in the NEON experiment for observing coherent neutrino-nucleus scattering during an engineering run at a nuclear reactor Choi et al. (2023). However, the design revealed some weaknesses, as liquid scintillator leakage into the detector caused a gradual decrease in crystal light yield and increase of PMT-induced noise.
Since then, We have improved the crystal encapsulation by separating it into two components: an inner structure to maintain a stable coupling between the PMTs and crystal, and an outer OFC case to prevent the infiltration of external air and liquid scintillator. These improvements were successfully applied to the NEON experiment, which has been collecting stable physics data for over two years Choi et al. (2024a).
II.1 Design of COSINE-100U Encapsulation
To maximize light collection efficiency, we used 3-inch diameter crystals to match the PMT diameters in the NEON experiment Choi et al. (2023). However, COSINE-100 crystals have larger diameters than the PMTs, as summarized in table 1. To address this, we designed the crystal edges to guide light from the larger crystal diameter to the 3-inch PMTs by machining beveled edges at a 45 angle, as shown in Fig. 2 (a) for the C6 crystal. During this process, the effective area of the 3-inch PMT photocathode, typically 72 mm, was considered, and the crystal was machined to a 70 mm diameter at the ends to optimize PMT attachment.
Figure 2: An example of the inner structure for the C6 crystal, designed to maintain a stable coupling between the PMTs and crystal. (a) The design of the C6 machining for guiding light to 3-inch PMTs. (b) A PTFE structure that covers the crystal and connects it to the PMTs. (c) An illustration of the crystal-PMT assembly, including 2 mm thick optical pad for coupling.
To couple the crystal with the PMTs, we designed a 5 mm-thick PTFE structure to surround the crystal, as shown in Fig. 2 (b). This PTFE structure is horizontally separable at the center, allowing the crystal to be placed inside. Each end of PTFE structure is connected to the PMT using a PTFE ring with brass bolts, applying pressure to the optical pad between the crystal and the PMT for optical light coupling, as illustrated in Fig. 2 (c).
The outer tube, made from OFC, follows the updated design of the NEON experiment Choi et al. (2024a), but with a larger diameter and length. As shown in Fig. 3 (a), the C6 tube has a length of 660 mm, a diameter of 140 mm, and a thickness of 2 mm. The side flanges (Fig. 3 (b)) are 20 mm-thick to provide sufficient pressure for a tight seal. Two grooves were machined into each side to accommodate PTFE gaskets, as shown in Fig. 3 (c). The two OFC lids (Fig. 3 (d)), each 20 mm thick, are attached to the tube flanges using twelve brass bolts through the PTFE gaskets. Each lid contains three holes for passing one high-voltage cable and two signal cables for the anode and dynode readouts Adhikari et al. (2018c). Waterproof cable glands secure the lids, preventing the ingress of liquid scintillator and external air through the cable exit holes. The final assembly of the inner and outer structures is illustrated in Fig. 3 (e).
Figure 3: An example of the outer structure for the C6 crystal, designed to prevent the infiltration of outside air and liquid scintillator. (a) The OFC tube for the C6. (b) Side view of the OFC tube showing the flange with two grooves for tight sealing using gaskets. (c) Design of the PTFE gasket. (d) OFC lid, which couples with the flange of the tube through the PTFE gasket. (e) Illustration of the combined inner and outer structure for C6
II.2 Crystal Machining and Encapsulation
To upgrade the crystals for the COSINE-100U experiment, the existing encapsulation was first removed, and the crystal edges were machined, as shown in Fig. 4. This process was performed by a specialized machining company. A dry room with a dedicated lathe machine was prepared for removing the original encapsulation and shaping the crystal edges to improve light guidance. During machining, mineral oil was continuously poured over the crystal to prevent cracks and suppress NaI(Tl) dust in the environment. Given the highly hygroscopic nature of NaI(Tl) crystals, the machined crystals were stored in a dry storage box and quickly transferred to a low-humidity, N2 gas-flushed glovebox to avoid exposure to atmospheric moisture.
Figure 4: Machining and delivery process of the C1 crystal. (a) The COSINE-100 C1 bare crystal after the removal of the original encapsulation. (b) Machining of C1 using a lathe machine. (c) The crystal storage box to transport the machined crystal to glovebox. The storage box was filled with dehumidifiers and flushed with N2 gas to prevent moisture exposure.
The humidity level inside the glovebox was maintained at below a few tens of ppm (H2O) using a molecular sieve trap and N2 gas flushing. Before polishing the crystals, N2 gas was flushed at a high flow rate of 15 liters per minute for 2 hours, which was then reduced to 5 liters per minute to minimize radon levels inside the glovebox.
After machining, the crystals were contaminated with mineral oil, NaI(Tl) powder, and other debris, as shown in Fig. 5 (a). To clean the crystal surfaces, they were gently wiped with anhydrous ethanol. Due to the small amount of water in the ethanol, this process slightly removed the outer crystal layer. Each end of the crystal was then mirror-polished using a polishing pad and SiO2 abrasives in two-steps: first with 3 $\mu$m and then with 0.5 $\mu$m particle sizes. After polishing, all surfaces were wiped with cleanroom wipes soaked in anhydrous isopropanol, resulting in a shiny finish on the crystal surfaces, as shown in Fig. 5 (b). The stability of the polished crystal surfaces was tested in the glovebox for a week, during which no visible changes were observed.
Figure 5: Example of the COSINE-100U crystal encapsulation process. (a) The C1 crystal was delivered to a low-humidity glovebox after machining for light guidance. (b) The crystal surface was cleaned and polished. (c) The inner structure was assembled by directly attaching the PMTs to the crystal using an optical pad. (d) The inner structure was inserted into the outer OFC tubes and sealed with two OFC lids using PTFE gaskets. (e) The outer lids feature three outlets for signal and high-voltage cables, secured with waterproof cable glands.
[…]
The polished crystals were wrapped with 250 $\mu$m thick Teflon sheets to enhance light collection. They were directly attached to the PMTs using 2 mm-thick optical pads, held in place by a PTFE inner structure with six screw holes, as shown in Fig. 5 (c). PTFE pressure rings were used to apply sufficient pressure between the crystal and PMTs for proper optical coupling, secured with brass bolts, nuts, and washers. Special locking washers were used to prevent loosening due to vibrations during transport and operation. The entire assembly was then placed inside a copper case and sealed with copper lids using brass bolts and nuts, as shown in Fig. 5 (d). To prevent external contamination, PTFE gaskets and double layers of PTFE reflective sheets were inserted at the flange joints. Cables were routed through cable glands at the center of the copper lid, as shown in Fig. 5 (e).
III Sea Level Measurements
III.1 Measurement Setup
Upon assembling the COSINE-100 crystals, the Yemilab facility was not yet ready for the operation of the COSINE-100U experiment. For the initial crystal characterization, we employed a simple shielding setup at sea level in the experimental hall of the Institute for Basic Science (IBS) in Korea. This setup consisted of two layers of shielding: 10 cm-thick lead and 20 cm-thick liquid scintillator, which also functioned as an active veto detector. The liquid scintillator was housed within a 124.5 cm $\times$ 49.5 cm $\times$ 49.5 cm cubic stainless steel box. Three 8-inch PMTs were used to read the signals from the liquid scintillator. This setup was initially developed as a prototype detector for the NEOS experiment Ko et al. (2017) and was reused for this test. Inside the container, an acrylic table was used to install one NaI(Tl) crystal for the first characterization, as shown in Fig. 6.
Figure 6: (a) Schematic view of the shield structure used in the sea level measurement at the Institute for Basic Science. (b) The C1 crystal is installed in this setup for light yield and stability measurements.
III.2 Light Yield Measurement
We first measured the light yield of the newly encapsulated crystal by irradiating it with 59.54 keV $\gamma$-rays from a 241Am source. The mean charge corresponding to a single photoelectron (SPE) was determined by analyzing trailing isolated cluster pulses in the 5–7 and 6–8 $\mu$s windows, which were 2.6–4.6 and 3.6–5.6 $\mu$s from the typical pulse start Adhikari et al. (2018c), to avoid multiple photoelectrons. We simultaneously fit the 5–7 and 6–8 $\mu$s clusters with models of up to four photoelectrons clusters Choi et al. (2024b, a), as shown in Fig. 7. The number of photoelectrons (NPE) was calculated by dividing the integrated charge of the main pulse (within 5 $\mu$s of the pulse start) by the charge of SPE.
Figure 7: The charge distribution of the isolated clusters in the 5–7 $\mu$s (red dots) and 6–8 $\mu$s (blue dots) windows. The distribution is modeled with up to four NPE clusters, simultaneously fitted for both the 5–7 and 6–8 $\mu$s windows, yielding a mean charge of 597.2 ADC for SPE.
Figure 8 (a) shows the spectra of 241Am calibration, representing NPEs from the COSINE-100 (black dashed line) and COSINE-100U (red solid line) encapsulation setups. The COSINE-100U setup provides approximately a 35% increase in light yield. From the mean of 59.54 keV peak, we observed a light yield of 19.6$\pm$0.3 NPE/keV for COSINE-100U C6, with similar improvements observed for all crystals, as summarized in Table 1.
Figure 8 (b) shows the 241Am calibration spectra in terms of energy. The COSINE-100U setup clearly demonstrates improved energy resolution compared to COSINE-100, benefiting from the increased light yield. Similar improvements were observed across all crystals.
Figure 8: Measured spectra from the C6 crystal in the original COSINE-100 encapsulation and the newly designed COSINE-100U encapsulation at 59.54 keV using a 241Am source. (a) Comparison of light yield between COSINE-100 (black dashed line) and COSINE-100U (red solid line). (b) Comparison of energy resolution between COSINE-100 (black dashed line) and COSINE-100U (black dashed line).
In the COSINE-100 experiment, C5 and C8 recorded relatively low light yields due to their 5-inch diameter optical windows, which were initially designed to accommodate 5-inch PMTs. This reduced light yield led to the exclusion of C5 and C8 from the low-energy dark matter search analysis. Additionally, C1 was excluded due to unexpected large noise events from its PMTs. Although the COSINE-100 experiment used eight crystals with a total mass of 106 kg, the effective detector mass used for the main physics analysis was only 61.4 kg Adhikari et al. (2021a, 2022a).
Upon disassembling C5 and C8 for the COSINE-100U encapsulation, we observed liquid scintillator leakage inside the crystals through thin Mylar windows, which were initially used for low-energy x-ray calibration, such as the 5.9 keV signal from 55Fe. This leakage created grooves a few millimeters deep in the crystal surfaces. During the polishing process, we smoothed these grooves, but some remained, resulting in a reduced light yield of approximately 16.5 NPE/keV for C5 and C8. While this yield is lower than that of other COSINE-100U crystals, it is higher than the light yields of COSINE-100’s good-quality detectors, as summarized in Table 1. As a result, these two crystals can now be used for the physics analysis in the COSINE-100U experiment, increasing the effective detector mass to 99.1 kg.
Table 1: Summary of the COSINE-100 and COSINE-100U crystals, including their dimensions, mass, and light yields. The total mass of the COSINE-100 crystals was 106.3 kg, but only 61.4 kg of crystals were used for the dark matter search. Due to the machining process, the total mass of the COSINE-100U is reduced to 99.1 kg, with all crystals available for dark matter search analysis. Overall, the COSINE-100U setup provides a significantly increased light yield.
Crystal Size (inches) Mass [kg] Light yield [PEs/keV]
diameter×$\times$×length COSINE-100 COSINE-100U COSINE-100 COSINE-100U
C1 5.0$\times$7.0 8.3 7.1 14.9 $\pm$ 1.5 22.4 $\pm$ 0.5
C2 4.2$\times$11.0 9.2 8.7 14.6 $\pm$ 1.5 20.1 $\pm$ 0.5
C3 4.2$\times$11.0 9.2 8.7 15.5 $\pm$ 1.6 20.4 $\pm$ 0.4
C4 5.0$\times$15.3 18.0 16.9 14.9 $\pm$ 1.5 20.7 $\pm$ 0.4
C5 5.0$\times$15.5 18.3 17.2 7.3 $\pm$ 0.7 16.8 $\pm$ 0.5
C6 4.8$\times$11.8 12.5 11.7 14.6 $\pm$ 1.5 19.6 $\pm$ 0.3
C7 4.8$\times$11.8 12.5 11.6 14.0 $\pm$ 1.4 20.2 $\pm$ 0.5
C8 5.0$\times$15.5 18.3 17.2 3.5 $\pm$ 0.3 16.2 $\pm$ 0.4
Total 106.3 (61.4) 99.1
III.3 Internal Background and Stability Measurement
After the 241Am measurements, each crystal underwent approximately two weeks of background measurements in the sea level shield setup, as shown in Fig. 6. Due to the relatively thin layers of lead and liquid scintillator, as well as the high muon flux at sea level, we could not achieve the low-background levels of the COSINE-100 experiment at Y2L. However, we were still able to study internal $\alpha$ background and assess the stability of the encapsulation.
Figure 9: Single-hit low-energy spectra of C6 at the COSINE-100 experiment (dashed line) and the sea level simple shield using COSINE-100U encapsulation (red-line)
Events coincident with the liquid scintillator with energies above 80 keV and within a 200 ns coincidence window were categorized as multiple-hit events, while all other events were categorized as single-hit events. Figure 9 shows the single-hit low-energy spectra of C6 from this measurement, using the upgraded COSINE-100U encapsulation (red-solid line) compared to the same crystal in the COSINE-100 experiment. As seen in the figure, the sea level measurement had significantly higher background rates due to reduced shielding and increased muon-related backgrounds. A peak around 33 keV was observed in this sea level measurement with the COSINE-100U encapsulation. This could be due to external contributions, with the K-shell dip of non-proportional scintillation light Lee et al. (2024a) possibly contributing to the peak. Additionally, x-rays from Ba in the PMT glass, as well as Cs and In in the photocathodes, may also contribute. The removal of the 12 mm-thick quartz layer between the crystal and PMTs in the new encapsulation may have enhanced these x-ray signals.
Figure 10: (a) Charge weighted mean decay time is shown for C6. Bulk $\alpha$ (red solid rectangle) and surface $\alpha$ (blue dashed rectangle) are indicated. (b) Energy spectra of $\alpha$ candidate events in this measurement (red-solid line) is compared with COSINE-100 setup (black-dashed line)
To measure internal $\alpha$ activity, we utilized charge-weighted mean decay time to distinguish between $\alpha$ and beta/gamma events, as shown in Fig. 10 (a), where $\alpha$s form a distinct cluster with shorter decay times, clearly separated from the beta/gamma events. The bulk $\alpha$ contamination from 210Po is highlighted in the red solid box, while low-energy surface $\alpha$ contamination Adhikari et al. (2024), with energy in the 1–2 MeV range, is indicated by the blue dashed box. Bulk $\alpha$ contamination originates from impurities introduced during the crystal growing process and is expected to be consistent with the COSINE-100 measurements, decreasing over time due to the decay of 210Pb ( with a half-life of 22.3 years). Surface $\alpha$ contamination, on the other hand, may occur on the crystal surface or on the PTFE reflective sheet during the encapsulation process. We used the event rate of 1–2 MeV $\alpha$-particles as an indicator of surface contamination, as shown in Fig. 10 (b). Our careful encapsulation process has minimized surface contamination. Table 2 summarizes internal $\alpha$ background measurements of the new COSINE-100U encapsulation compared to the COSINE-100 measurements near shutdown in March 2023. The bulk $\alpha$ measurements show a clear decrease in the COSINE-100U setup, consistent with the decay of internal 210Pb, while surface $\alpha$ rates are generally lower than in COSINE-100, though some crystals show slightly higher rates. We plan to systematically study surface $\alpha$ contamination by varying surface treatment methods using sample crystals to better understand the causes of contamination.
Table 2: The summary of COSINE-100U $\alpha$ measurements.
Crystal number Bulk alpha [mBq/kg] surface alpha [nBq/cm2]
COSINE-100 COSINE-100U COSINE-100 COSINE-100U
1 2.59 $\pm$ 0.01 2.49 $\pm$ 0.02 33.07 $\pm$ 3.70 45.22 $\pm$ 5.99
2 1.69 $\pm$ 0.01 1.63 $\pm$ 0.02 39.61 $\pm$ 3.76 30.42 $\pm$ 10.14
3 0.63 $\pm$ 0.01 0.59 $\pm$ 0.01 71.73 $\pm$ 5.06 30.89 $\pm$ 8.57
4 0.64 $\pm$ 0.01 0.60 $\pm$ 0.01 24.27 $\pm$ 2.31 51.48 $\pm$ 5.76
5 - 1.69 $\pm$ 0.02 - 99.45 $\pm$ 11.72
6 1.52 $\pm$ 0.01 1.42 $\pm$ 0.01 120.9 $\pm$ 5.90 39.11 $\pm$ 5.83
7 1.51 $\pm$ 0.01 1.43 $\pm$ 0.01 95.35 $\pm$ 5.24 61.70 $\pm$ 7.32
8 - 1.50 $\pm$ 0.02 - 24.71 $\pm$ 4.75
To verify the stability of the assembled crystals, each crystal was monitored in the sea level test measurement facility for approximately two weeks. Specifically, the 33 keV and the 46.5 keV radiation peaks were tracked to check the stability of the encapsulation. Figure 11 (a) shows the data collected over this period, plotted in 100-hour intervals, demonstrating no noticeable shifts in the peak positions. This indicates that the crystal-PMT coupling remained robust and that no infiltration of liquid scintillator or air occurred.
After completing the background measurements, the upgraded crystals were delivered to Yemilab to minimize cosmogenic activation. The crystals were stored in nitrogen-flushed clean storage. Only the 241Am source measurement, conducted inside a dark box, was used to monitor any variation in light yield, as shown in Fig. 11 (b). We observed consistent light yields from 59.54 keV peak, indicating stable conditions of the crystal encapsulation.
Figure 11: Stability of the C1 encapsulation. (a) Gain stability of sea level background measurements for approximately two-week long periods was shown with stable gain condition. (b) Measurements of 241Am source to 59.54 keV at sea level measurement (black dashed line) and Yemilab measurement (red solid line) are presented. With two-months period, no significant change of gain was observed.
IV Yemilab Preparation
IV.1 Decommissioning of COSINE-100
The COSINE-100 experiment, which operated at Y2L, concluded in March 2023 in preparation for the relocation of the experimental site to Yemilab Park et al. (2024); Kim and Lee (2024) and the detector upgrade for the COSINE-100U experiment. The decommissioning of the detector was completed by October 2023, as shown in Fig. 12, and all materials were delivered to Yemilab for the installation of COSINE-100U.
Figure 12: Photos taken during the decommissioning of the COSINE-100 detector at Y2L. (a) The external shield of the COSINE-100 before decommissioning. (b) Removal of outside electronics and muon detectors, with the decommissioning of lead bricks in progress. (c) All external passive shields were removed, and the PMTs for the liquid scintillator active veto detector were detached. (d) Complete removal of the COSINE-100 shield from Y2L.
IV.2 Yemilab Preparation
Yemilab is a newly constructed underground laboratory in Korea, completed in September 2022, located in Jeongseon, Gangwon Province, at a depth of 1,000 m Park et al. (2024); Kim and Lee (2024). The facility offers approximately 3,000 m2 of dedicated experimental space. The underground tunnel accommodates 17 independent experimental spaces, one of which is dedicated to the COSINE-100U experiment, as shown in Fig. 13 (a). The tunnel can be accessed via a human-riding elevator through a 600 m vertical shaft and then by electric car through a 780 m horizontal access tunnel with a 12% downward slope. The surrounding rock is primarily limestone. Ongoing radioactivity measurements of rock samples using ICP-MS and HPGe detectors show that the preliminary results are generally consistent with, or slightly lower than, those from Y2L.
Figure 13: (a) Schematic view of the underground experimental area at Yemilab and (b) The COSINE-100U fridge room to operate at –30 C.
[…]
Preliminary measurements of muon flux at Yemilab indicate a flux of 1.0×$\times 10^{-7}\mu$/cm2/s, which is four times lower than the muon flux at Y2L, measured at 3.8×$\times 10^{-7}\mu$/cm2/s Prihtiadi et al. (2018). Overall, the background environments at Yemilab are significantly better than that of Y2L, leading to reduced external radioactive background contributions.
We have prepared a warehouse-type refrigerator to serve as the COSINE-100U detector room, as shown in Fig.  13 (b). The plan is to operate the COSINE-100U experiment at –30 C to enhance light yield and improve pulse shape discrimination for nuclear recoil events Lee et al. (2022). The COSINE-100U detector room measures 4 m in width, 6 m in length, and 4 m in height, and is located at the front of the COSINE tunnel. A 10 kW cryocooler will maintain the room temperature.
IV.3 Shielding Installation
Figure 14: Shield installation of COSINE-100U at Yemilab. (a) The liquid scintillator active veto detector, composed of a 1 cm thick acrylic box surrounded by 3 cm think OFC, was installed with 5-inch PMTs for readout. (b) A 20 cm thick lead layer surrounds the copper container for the liquid scintillator. (c) Inside of the acrylic box for the liquid scintillator, the C5 and C8 were installed.
Inside the COSINE-100U fridge room, we installed shielding for the experiment. Most elements of the COSINE-100 experimental shield were recycled for the COSINE-100U setup to prevent external radiation from various sources and provide an active veto for internal or external contamination Adhikari et al. (2018a). This shield consists of a 4-layer nested arrangement of components, starting from the inside: 40 cm liquid scintillator, 3 cm of copper, 20 cm of lead, and 3 cm of plastic scintillator. The liquid scintillator Adhikari et al. (2021b) and plastic scintillator Prihtiadi et al. (2018) layers actively tag radioactivity from internal contamination, external radiation, and muon events.
The COSINE-100 shield utilized a steel skeleton to support heavy elements and allow access to the inner structure with a mechanical opening system Adhikari et al. (2018a). However, this design inherently included approximately 4 tons of steel inside the lead shields. In contrast, the COSINE-100U shield does not use a steel skeleton. Instead, heavy materials, such as lead bricks, are stacked on a precisely leveled steel plate, similar to the shield used in the NEON experiment Choi et al. (2023), as shown in Fig. 14. To reinforce the top structure, 5 cm×$\times$10 cm square stainless steel pipes, each 180 cm long, will support the lead bricks. Figure 15 shows the detector geometry of the experimental setup.
We produced 2,400 L Linear Alkyl-Benzene(LAB) based liquid scintillator, following a recipe similar to that used in the COSINE-100 experiment Adhikari et al. (2018a, 2021b). The old COSINE-100 liquid scintillator will be repurposed for test measurements facilities at Yemilab.
IV.4 Physics Operation Plan
For the operation of the COSINE-100U, all electronics, including preamplifiers, flash analog-to-digital converters, high voltage power supplies, and the computer server for data acquisition, will be installed in a –30C environment. While the stability of individual components had already been tested at this low temperature, integrated tests at Yemilab are planned for a few weeks. During this test, only two crystals, C5 and C8, as shown in Fig. 14 (c), will be installed, and the liquid scintillator will not yet be filled. We have just started room temperature measurements to obtain reference data for this test.
If no critical issues arise from the NaI(Tl) crystal detectors, electronics, or data acquisition system, we will proceed with the installation of the remained detector components, including the other six crystals, the liquid scintillator, and the top lead bricks and outer muon plastic scintillator panels. This entire process is expected to be completed with a month. If everything proceeds smoothly, we anticipate that the COSINE-100U experiment will begin physics operations in October 2024.
V Expected Background
We have gained a precise understanding of the backgrounds in the COSINE-100 detector through Geant4-based simulations Adhikari et al. (2018d, 2021c); Yu et al. (2024a). To account for COSINE-100U-specific background contributions, we constructed detector geometry for use in the Geant4-based simulation, as shown in Fig. 15. Since the COSINE-100U experiment uses the same crystals as COSINE-100, with only minor machining and surface polishing, we expect the majority of background contributions in COSINE-100U, particularly from internal contaminants, to be very similar to those observed in the COSINE-100 experiment.
Figure 15: COSINE-100U detector geometry used in the Geant4 simulation: (a) bird’s-eye view and (b) front view. Eight NaI(Tl) crystals are supported by the acrylic table inside the liquid scintillator. Nine 5-inch PMTs in each side are attached to read photons from the liquid scintillator. The outer lead shield and plastic scintillators are also installed. To support the shielding on the top, stainless steel support pipes are installed. The entire shieling setup is placed on a precisely labeled steel base.
However, a few differences are expected due to the redesigned crystal encapsulation. The encapsulation components were replaced, and an additional inner PTFE structure was incorporated. We measured the radioactivity levels of all encapsulation components, as summarized in table 3. As we carefully selected all materials, the contamination levels of the new encapsulation materials are much lower than those of the PMT and PMT base. Based on our understanding of the COSINE-100 backgrounds Yu et al. (2024a) and the measured contamination levels of the encapsulation materials, we simulated the expected background contributions. As in COSINE-100, the PMTs remain the dominant source of external background contamination.
Table 3: Measured radioactivity contaminants in detector components inside the shielding. The radioactivities were measured using high-purity germanium detectors at Y2L; upper limits are quoted with a 90% confidence level. The PMTs and PMT bases are measured in units of mBq/unit, while the other components are measured in units of mBq/kg.
Material 238U 232Th 40K
Copper 0.27 ± 0.07 0.25 ± 0.07 $<$ 1.24
PTFE $<$ 0.19 $<$ 0.38 8.88 ± 1.32
Brass bolts $<$ 0.7 $<$ 0.7 $<$ 6.1
PMTs 60 ± 10 12 ± 5 58 ± 5
PMT Base 12.2 ± 2.5 1.7 ± 0.4 6.2 ± 2.5
The polishing of all crystal surfaces and the replacement of the Teflon lapping films may result in different surface contamination levels in the COSINE-100U crystals. Generally, we observed fewer $\alpha$ particles with partial energy deposition (1–2 MeV measured energy), which may suggest lower surface contamination. However, for this study, we conservatively assume the same surface contamination background contributions as observed in the COSINE-100 experiment.
Although the same PMTs are used, the removal of the 12 mm quartz layer could potentially increase background contributions from the PMTs to the crystals. We simulated these background contributions in the COSINE-100U geometry (Fig. 15), assuming the same contamination levels as in the COSINE-100 experiment Yu et al. (2024a). The absence of the 12 mm quartz shield may enhance the x-ray contribution from the PMTs, but this effect is primarily observed at energies above 20 keV, with no significant differences in the signal region below 6 keV, as shown in Fig. 16.
Figure 16: Expected energy spectra of the C6 crystal in the COSINE-100U setup compared with the COSINE-100 background Yu et al. (2024a). (a) Low-energy spectra from the anode readout of COSINE-100 (black solid line) compared with the expected background of the COSINE-100U setup (red solid line). The main differences in PMT contributions are separated as dashed lines. (b) High-energy spectra from the dynode readout are presented in the same manner. Although an increase in background contributions from the PMTs is expected in the COSINE-100U encapsulation, its impact on the signal region below 6 keV is negligible.
Figure 16 shows the expected background levels of the COSINE-100U C6 crystal compared with those of the COSINE-100 setup, based on the assumptions discussed above and the Geant4-based simulation. The main differences are caused by PMT x-rays resulting from the absence of the 12 mm quartz windows; however, no significant difference is observed in the low-energy signal regions.
VI Sensitivity of the COSINE-100U experiment
With the improved performance of higher light yields in the COSINE-100U detectors, along with background levels similar to those observed in the COSINE-100 experiment in the low-energy signal region, we evaluate the sensitivity of the COSINE-100U experiment for detecting dark matter, particularly for spin-dependent WIMP-proton interaction. We assume one year of operation, using the measured light yields at room temperature as summarized in table 1, and the expected background levels discussed in Section V, based on the COSINE-100 measurement Yu et al. (2024a). While the current COSINE-100 data analysis has reached an 8 NPE threshold Yu et al. (2024b), further improvements using machine learning techniques and simulated waveform of NaI(Tl) crystals Choi et al. (2024b) are expected to lower the threshold to 5 NPE, a level already achieved by the COHERENT experiment with CsI(Tl) crystal Akimov et al. (2017). For the sensitivity analysis of the COSINE-100U experiment, we assume a 5 NPE analysis threshold for each crystal.
We generate WIMP interaction signals with and without the Midgal effect Migdal (1941); Ibe et al. (2018); Adhikari et al. (2022b). These signals are simulated for various interactions and masses within the standard WIMP galactic halo model Lewin and Smith (1996); Freese et al. (2013). Form factors and proton spin values of the nuclei are implemented using the publicly available dmdd package Gluscevic and McDermott (2015); Gluscevic et al. (2015); Anand et al. (2014); Fitzpatrick et al. (2013); Gresham and Zurek (2014). The electron-equivalent energy of the nuclear recoil is reduced using nuclear recoil quenching factors, which represent the ratio of scintillation light yield from sodium or iodine recoil relative to that from electron recoil for the same energy. Recently measured quenching factor values Lee et al. (2024b) were used, and the inclusion of the Migdal effect in the NaI(Tl) crystals follows our previous study Adhikari et al. (2022b). The measured electron-equivalent energy is then converted to NPE based on the light yield shown in Table 1.
Poisson fluctuations in the measured NPE are considered for detector resolution, using a recently developed waveform simulation package Choi et al. (2024b), which has already provided a good description of energy resolution in the low-energy signal region of COSINE-100 data Lee et al. (2024a). We use an ensemble of simulated experiments to estimate the sensitivity of the COSINE-100U experiment, expressed as the expected cross-section limits for the WIMP-proton spin-dependent interactions in the absence of signals. For each experiment, a simulated spectrum is generated under a background-only hypothesis based on assumed background levels. Gaussian fluctuations of background components from the COSINE-100 measurement Yu et al. (2024a), along with COSINE-100U-specific background contributions discussed in Section V, and Poisson fluctuations in each energy bin are incorporated into each simulated experiment.
[…]
We then fit the simulated data with a signal-plus-background hypothesis, applying flat priors for the signal and Gaussian constraints for the backgrounds. Systematic uncertainties affecting the background model are included as nuisance parameters Adhikari et al. (2021a). A Bayesian approach is used to analyze the single-hit energy spectrum between 5 NPE and 130 NPE for each WIMP model, covering several WIMP masses. Marginalization is performed to obtain the posterior probability density function for each simulated sample, allowing us to set the 90% confidence level exclusion limits. The 1,000 simulated experiments result in 68% and 95% bands of the expected limits, as shown in Fig. 17.
The expected limits from the COSINE-100U experiment are compared with the current best limits on low-mass WIMP-proton spin-dependent interactions from PICO-60 Amole et al. (2019), CRESST-III Li Angloher et al. (2022), NEWS-G Arora et al. (2024), and Collar Collar (2018). Leveraging the odd-proton numbers and relatively small atomic mass of sodium, the projected sensitivity of the COSINE-100U experiment can probe unexplored parameter spaces for WIMP masses below 4 GeV/c2, potentially reaching extremely low-mass regions as low as 20 MeV/c2 when considering the Migdal effect.
VII Summary
The COSINE-100U experiment represents a major upgrade from the COSINE-100 experiment, aimed at improving sensitivity to low-mass dark matter detection. After 6.4 years of successful operation at the Yangyang Underground Laboratory, the experiment was relocated to the newly constructed Yemilab in Korea, which provides a deeper underground environment with enhanced shielding from cosmic muons. Key improvements in the COSINE-100U experiment include the implementation of a novel crystal encapsulation technique that increases light collection efficiency by approximately 35%. We have evaluated the expected sensitivities of the COSINE-100U experiment, assuming a total mass of 99.1 kg, a 1-year operation period, and a 5 NPE energy threshold. Under these conditions, the COSINE-100U detector has the potential to explore previously uncharted parameter spaces for spin-dependent WIMP-proton interactions.
Acknowledgements.
We thank the Yemilab operation team for their dedicated support at Yemilab and the IBS Research Solution Center (RSC) for providing high performance computing resources. This work was supported by the Institute for Basic Science (IBS) under project code IBS-R016-A1, NRF-2021R1A2C3010989, NRF- 2021R1A2C1013761 and RS-2024-00356960, Republic of Korea; NSF Grants No. PHY-1913742, DGE-1122492, WIPAC, the Wisconsin Alumni Research Foundation, United States.